OpenStudy (anonymous):

The average salary for computer programmmers is $7,690 less than twice the average salary for carpenters. Combined, their avearge salaries are $99,830. Determine the average salary for each job.

OpenStudy (anonymous):

Let x = the salary of a carpenter:\[99830 = x + (2x - 7690)\]\[99830 = 3x - 7690\]\[107520 = 3x\]\[35840 = x\] So a computer programmer's would be (2x - 7690), or 63990

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!

OpenStudy (anonymous):

No problem, click "Good Answer" next to my name to give me a medal :)